Name the property that is illustrated by statement: 2a+(3b+c)=(3b+c)+2a.
Simplify (-14+8)-(4-11).
Graph the equation 2x-y+1.
Find the slope of the line containing points (-4,3)and2,-5).

Answer by stanbon(48569)

Name the property that is illustrated by statement: 2a+(3b+c)=(3b+c)+2a.
Associative Property
--------------------------------
Simplify (-14+8)-(4-11).
-14+8-4+11
-18+19
1
-----------------------------
Graph the equation 2x-y+1.
You don't have an equation.
----------------------------------
Find the slope of the line containing points (-4,3)and2,-5).
slope = (-5-3)/(2--5)=-8/7
---------------------------------------------------------------
Cheers,
Stan H.